OpenStudy (anonymous):

Write 2[ln x − ln(x + 1) − ln(x − 1)] as the logarithm of a simple quantity

OpenStudy (mertsj):

\[2\ln [ x-[\ln (x+1)+\ln (x-1)]]=2\ln [\frac{ x}{(x+1)(x-1)}]=\ln \frac{x}{(x+1)(x-1)}]^2\]

OpenStudy (mertsj):

Last part got cut off. \[\ln [\frac{x}{(x+1)(x-1)}]^{2}\]

OpenStudy (accessdenied):

A few neat properties of logarithms to know about... ln(a) - ln(b) = ln( a/b ) a * ln(b) = ln(b^a) 2 [ ln(x) - ln(x + 1) - ln(x - 1) ] Let's work with what is inside of those square brackets first and we'll replace it later. ln(x) - ln(x + 1) - ln(x - 1) I'll rewrite with first two terms in parentheses (its not needed, I do it to show process.) = ( ln(x) - ln(x + 1) ) - ln(x - 1) Here, we can use that first property. = ln(x / (x+1)) - ln(x - 1) And then again, use that property. = ln( (x / (x + 1)) / (x - 1) ) a/b = a*(1/b) = ln( (x / (x + 1)) * (1 / (x - 1)) ) We can see that (x + 1)(x -1) will become x^2 - 1 = ln( x / (x^2 - 1) ) Now, we'll replace that stuff in the brackets with what we found. 2[ ln( x / (x^2 - 1) ) ] Here's where the second property listed comes in. ln( (x / (x^2 - 1))^2 )